OpenStudy (anonymous):

A tugboat goes 180 miles upstream in 15 hours. The return trip downstream takes 5 hours. Find the speed of the tugboat without a current and the speed of the current. The speed of the tugboat is ?mph, and the speed of the current is ? mph. Simplify your answers

OpenStudy (anonymous):

36 and 24?

hero (hero):

Hint use r = speed of boat w = speed of water d = (r - w)t for upstream d = (r + w)t for downstream Since we know the distance is 180 miles and we know it takes 15 hours upstream and 5 hours downstream then set up the following equation: 180 = 15(r - w) 180 = 5(r + w) Afterwards, that will give you a system of two equations, which you will be able to solve for r and w.

OpenStudy (anonymous):

is it 36 and 24?

hero (hero):

For some reason I get something different.

OpenStudy (anonymous):

You are prob right. Im not sure of my answer

hero (hero):

Show your steps.

OpenStudy (anonymous):

I think it is 36 and 12

OpenStudy (anonymous):

Is taht what you got hero?

hero (hero):

@Zeph, show your steps.

hero (hero):

Don't just post answers without explaining yourself.

OpenStudy (anonymous):

I divided 180 by 15mph. gave me 12. so it was going 12 mph upstream. Then i divided 180 by 5 and came up with 36 so i subtracted 12 from that and got my answer

OpenStudy (anonymous):

For the first equation I divided 180 by 15 to get 12. The second equation I divided 180 by 5 and got 36.

hero (hero):

Strange. That is not consistent with the equations I posted above.

OpenStudy (anonymous):

To be honest your equations confused me more

hero (hero):

You're familiar with d = rt right?

OpenStudy (anonymous):

distance and round trip?

OpenStudy (anonymous):

rate and time

OpenStudy (anonymous):

distance rate and time

hero (hero):

distance = rate times time C'mon, you have to know that. It's a basic equation.

OpenStudy (anonymous):

Im just flusterd sorry

hero (hero):

Okay, so if we add the upstream and downstream part to it you get the equations I wrote above.

hero (hero):

- w = upstream or against the current + w = downstream or with the current

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

15 hours at 12 mph

OpenStudy (anonymous):

36 mph downstream

hero (hero):

Solving for upstream and downstream: 180 = 15(r - w) 180 = 5(r + w) 180/15 = r - w 180/5 = r + w 12 = r - w 36 = r + w Add the above to get 48 = 2r 48/2 = r 24 = r

hero (hero):

r = speed of the boat

hero (hero):

Now using one of the equations above, solve for w: 36 = r + w 36 = 24 + w 36 - 24 = w 12 = w

hero (hero):

So I get: 24mph for the speed of the boat 12mph for the speed of the water

OpenStudy (anonymous):

so where did i screw up?

hero (hero):

By believing that you can solve it by simply dividing 180/15 and 180/5

OpenStudy (anonymous):

ahhhhhhh good point